OpenStudy (anonymous):

oh no, she's back, LOL! ok, just a quick question for you folks. x^2 + 2xy + y^2 with x=2 and y= -3. I get the answer 1. Is this correct and will you show me by equation how you got it? I just want to make sure I do the rest of the problems correctly. I found the Order of Operations was necessary.

OpenStudy (anonymous):

when u plug values x=2 and y=-3 in the eqn x^2+2xy+y^2 u get (2)^2+2*2*(-3)+(-3)^2 =4-12+9 =13-12 =1

OpenStudy (anonymous):

plug in your x for the x's and y's for the y's so your problem looks like this 2^2 + 2*2*-3 + -3^2 then you got 4-12-9 9-9 your answer is 0

Directrix (directrix):

-3^2 should be (-3)^2 which is 9.

OpenStudy (anonymous):

crap ur right im sorry im tired

OpenStudy (anonymous):

In the book the answer is 1 but that only helps me knowing, if I can figure out how they got it. If I order the terms at the last part to -12 + 4 -3^2 I do get the answer 1 like the book.

OpenStudy (anonymous):

it's ok Direct;) it's allowed to be worn out and tired;)

OpenStudy (anonymous):

Oh sorry Dawn, you are the one that's tired, well it's ok girlfriend;)

OpenStudy (anonymous):

thanks sorry for giving you the wrong answer is just slipped me that it should be positive not neg and that makes all the dif in the world in math

OpenStudy (anonymous):

LOL;) Thank you so much, I need to learn how to work the equation more than the answer because if I work it right, I get the answer:) Thank you again, we got it now:)

OpenStudy (anonymous):

I love this board, so many nice people that want to help. I wish I'd found you in my first term of Math:) Take care and see you later, Denise